OpenStudy (anonymous):

Can some one help solve... 3 * 18t^2 when t=1/3

OpenStudy (anonymous):

Just plug in 1/3 for t..

OpenStudy (anonymous):

3* 18* (1/3)^2 = 3*18*(1/9) = 3*2 = ??

OpenStudy (anonymous):

Not sure if i typed that right the ^ symbol is for raised to? Yes?

OpenStudy (anonymous):

That symbol is for raised to.. Or you can say that is the indication for Power or exponent..

OpenStudy (anonymous):

Yes. Raised to... like a power. This was seen to @waterineyes as\[(3)(18)t ^{2}\]where \[t=\frac{1}{3}\]

OpenStudy (anonymous):

My equation editor is not working this time that is why I can't show you more clearly..

OpenStudy (anonymous):

It's definitely faster if done correctly. :)

OpenStudy (anonymous):

I'm sooooo slow with the editor.

OpenStudy (anonymous):